OpenStudy (anonymous):

3(y – 5) + 2 = 5 solve for y please

jimthompson5910 (jim_thompson5910):

First distribute. Then combine like terms 3(y – 5) + 2 = 5 3*y – 3*5 + 2 = 5 3y – 15 + 2 = 5 3y – 13 = 5 What's next?

OpenStudy (anonymous):

y=6